OpenStudy (anonymous):

3 individuals take the elevator in 15-story building. What are the odds they all will get off on the same floor? Notice: they are all strangers to one another.

hartnn (hartnn):

can u find the probability first ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

@hartnn can u do it?

hartnn (hartnn):

the probability that 1 person gets off on 'n' floor = 1/15 so the probability that all 3 get off on same 'n' floor will be (1/15)(1/15)(1/15) git this ?

hartnn (hartnn):

*got this ?

hartnn (hartnn):

(1/15)(1/15)(1/15) = 1/3375 take p= 1/3375 calculate odds as p/(1-p)

OpenStudy (unklerhaukus):

isn't the probability that 1 person gets off on 'floor n' certain

OpenStudy (unklerhaukus):

there is only two degrees of freedom

OpenStudy (anonymous):

thanks @hartnn

OpenStudy (anonymous):

@hartnn can u please tell me the final answer?

hartnn (hartnn):

(1/3375)/(3374/3375) = 1/3374 = 2.96 *10^-4

OpenStudy (anonymous):

sorry but these are not the options of mine

hartnn (hartnn):

O.o

OpenStudy (anonymous):

options are 1/196 1/3375 1/2744 3/225

hartnn (hartnn):

if they are asking for chances(they might have used the word odds for chances or probability) then its 1/3375

OpenStudy (anonymous):

i also think so ... thanks for ur help

hartnn (hartnn):

welcome ^_^

OpenStudy (unklerhaukus):

the probability the first person gets off at stop n is 1 the probability that two people get off together at stop n is 1/15 the probability that three people get off together at stop n is 1/15 x 1/15 =1/225